OpenStudy (jamanorthi):

A 10ft fence is 8in long on a scale drawing. What's is the scale????????

OpenStudy (mertsj):

since 10 ft = 120 in I guess the scale is 8/120=1/15 so 1 inch on the drawing = 15 in in real life.

OpenStudy (jamanorthi):

I don't get it

OpenStudy (mertsj):

How many inches in 1 foot?